# Masked Generative Priors Improve World Models Sequence Modelling Capabilities

### v5

@@ Abstract
Deep Reinforcement Learning (RL) has become the leading approach for creating artificial agents in complex environments. Model-based approaches, which are RL methods with world models that predict environment dynamics, are among the most promising directions for improving data efficiency, forming a critical step toward bridging the gap between research and real-world deployment. In particular, world models enhance sample efficiency by learning in imagination, which involves training a generative sequence model of the environment in a self-supervised manner. Recently, Masked Generative Modelling has emerged as a more efficient and superior inductive bias for modelling and generating token sequences. Building on the Efficient Stochastic Transformer-based World Models (STORM) architecture, we replace the traditional MLP prior with a Masked Generative Prior (e.g., MaskGIT Prior) and introduce GIT-STORM. We evaluate our model on two downstream tasks: reinforcement learning and video prediction. GIT-STORM demonstrates substantial performance gains in RL tasks on the Atari 100k benchmark. Moreover, we apply Categorical Transformer-based World Models to continuous action environments for the first time, addressing a significant gap in prior research. To achieve this, we employ a state mixer function that integrates latent state representations with actions, enabling our model to handle continuous control tasks. We validate this approach through qualitative and quantitative analyses on the DeepMind Control Suite, showcasing the effectiveness of Transformer-based World Models in this new domain. Our results highlight the versatility and efficacy of the MaskGIT dynamics prior, paving the way for more accurate world models and effective RL policies. Here you can find our GitHub repository: https://github.com/Cmeo97/GitStorm/.
@@ 2.2 Masked Modelling for Visual Representations and Generation
The most relevant to our work is MaskGIT (Chang et al., 2022), a non-autoregressive decoding approach that consists of a bidirectional transformer model, trained by learning to predict randomly masked visual tokens. By leveraging a bidirectional transformer (Devlin et al., 2018), it can better capture the global context across tokens during the sampling process. Furthermore, training on masked token prediction enables efficient, high-quality sampling at a significantly lower cost than autoregressive models. MaskGIT achieves state-of-the art performance on ImageNet dataset and achieves a 64$\times$ speed-up on autoregressive decoding. The MaskGIT architecture has been applied to various tasks, such as video generation (Yan et al., 2023; Yu et al., 2023a; b) and multimodal generation (Mizrahi et al., 2024). For example, Yan et al. (2023) proposes TECO, a latent dynamics video prediction model that uses MaskGIT to model the prior for predicting the next timestep discrete representations, enhancing the sequence modelling of a backbone autoregressive transformer. Inspired by TECO, we adopt the use of MaskGIT prior for the world model, enhancing the sequence modelling capabilities, crucial for enabling and improving the agent policy learning behavior.
@@ 3 Method
Following DreamerV3 (Hafner et al., 2023) and STORM (Zhang et al., 2023), we define our framework as a partially observable Markov decision process (POMDP) with discrete timesteps, $t\in\mathbb{N}$, scalar rewards, $r_{t}\in\mathbb{R}$, image observations, $o_{t}\in\mathbb{R}^{h\times w\times c}$, and discrete actions. $a_{t}\in\{1,\ldots,m_{a}\}$. These actions are governed by a policy, $a_{t}\sim\pi(a_{t}\mid o_{1:t},a_{1:t-1})$, where $o_{1:t}$ and $a_{1:t-1}$ represent the previous observations and actions up to timesteps $t$ and $t-1$, respectively. The termination of each episode is represented by a Boolean variable, $c_{t}\in\{0,1\}$. The goal is to learn an optimal policy, $\pi$, that maximizes the expected total discounted rewards, $\mathbb{E}_{\pi}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r_{t}\right]$, where $\gamma\in[0,1]$ serves as the discount factor. The learning process involves two parallel iterative phases: learning the observation and dynamics modules (World Model) and optimizing the policy (Agent).
@@ 3.2 Dynamics Prior Head: MaskGIT Prior
[…]
@@
During inference, since MaskGIT has been trained to model both unconditional and conditional probabilities, we can sample any subset of tokens per sampling iteration. Following Yan et al. (2023), we adopt the Draft-and-Revise decoding scheme introduced by Lee et al. (2022) to predict the next latent state (Algorithm 1 and 2). During the draft phase, we initialize a partition $\bm{\Pi}$ which contains $T_{\text{draft}}$ disjointed mask vectors $\mathbf{m}$ of size (latent $(\text{latent dim}\div T_{\text{draft}})$, which together mask the whole latent representation. Iterating through all mask vectors in $\bm{\Pi}$, the resulting masked representations are concatenated with the hidden states $h_{t}$ from Eq. 1 and fed to the MaskGIT prior head that computes the logits of the tokens correspondent to $h_{t}$ and $\mathbf{m}^{i}$. Such logits are then used to sample the new tokens that replace the positions masked by $\mathbf{m}^{i}$. During the revise phase, the whole procedure is repeated $\Gamma$ times. As a result, when sampling the new tokens, the whole representation is taken into account, resulting in a more consistent and meaningful sampled state.
@@ 4.2 Results on Discrete Action Environments: Atari 100k
Figure 3 (Left) compares PI against the baselines. Noticeably, GIT-STORM presents $PI>0.5$ for all baselines, which indicates that, from a probabilistic perspective GIT-STORM would outperform each baseline on a random task Y from Atari 100k with a probability greater than $0.5$.  Figure 6 illustrates the Optimality Gap, while  Figure 7 presents the fraction of runs with score $>\tau$ for different human normalized scores; both confirm the trends observed so far. Moreover, a closer look to  Table 5 reveals that GIT-STORM presents an optimality gap of $0.500$, marginally beating DreamerV3, which reports $0.503$ and significantly outperforming all other baselines.
@@ A.1 Observation Module
[…]
@@ A.2 Policy Learning
[…]
@@
$\mathcal{L}_{\theta}=\frac{1}{BL}\sum_{n=1}^{B}\sum_{l=1}^{L}\left[-\text{sg}\left(\frac{G_{l}^{\lambda}-V_{\psi}(s_{l})}{\max(1,S)}\right)\ln\pi_{\theta}(a_{l}|s_{l})-\eta H(\pi_{\theta}(a_{l}|s_{l}))\right],\$ (12)
@@ A.3 Draft and Revise decoding scheme
6:$\bm{z}^{0}\leftarrow\textsc{MaskGIT Head}(\bm{z}^{\text{empty}},\bm{\Pi},h_{t};\theta)$
@@
[…]
@@
[…]
@@ Appendix C Full Results on RL Task
Table 5: Evaluation on the $26$ games in the Atari $100$k benchmark. We report mean scores as well as aggregated human normalized mean and median, Interquantile Mean (IQM), and Optimality Gap. Following the conventions of Hafner et al. (2021), scores that are the highest or within $5\%$ of the highest score are highlighted in bold.
@@ Appendix D Training Curves
Figure 8: Training profiles across all the checkpoints for the Atari 100k benchmark. The solid line represents the average over $5$ seeds while the fill area is defined in terms of maximum and minimum values corresponding to each checkpoint.
Figure 9: Performance evaluation across all the checkpoints for the DMC benchmark for GIT-STORM and STORM. The solid line represents the average over $5$ seeds while the fill area is defined in terms of standard deviation values corresponding to each checkpoint.
@@ Appendix H Hyperparameters
Table 11: Action mixer $\zeta_{t}=g_{\theta}(z_{t},a_{t})$. Concatenate denotes combining the last dimension of two tensors and merging them into one new tensor. The variable $A$ represents the action dimension, which ranges from $3$ to $18$ across different games. $D$ denotes the feature dimension of the Transformer. LN is an abbreviation for layer normalization (Ba et al., 2016).
@@
Random sample ($z_{t}$), Action ($a_{t}$) $32\times 32$, $A$

### v6

@@ Abstract
Deep Reinforcement Learning (RL) has become the leading approach for creating artificial agents in complex environments. Model-based approaches, which are RL methods with world models that predict environment dynamics, are among the most promising directions for improving data efficiency, forming a critical step toward bridging the gap between research and real-world deployment. In particular, world models enhance sample efficiency by learning in imagination, which involves training a generative sequence model of the environment in a self-supervised manner. Recently, Masked Generative Modelling has emerged as a more efficient and superior inductive bias for modelling and generating token sequences. Building on the Efficient Stochastic Transformer-based World Models (STORM) architecture, we replace the traditional MLP prior with a Masked Generative Prior (e.g., MaskGIT Prior) and introduce GIT-STORM. We evaluate our model on two downstream tasks: reinforcement learning and video prediction. GIT-STORM demonstrates substantial performance gains in RL tasks on the Atari 100k benchmark. Moreover, we apply Categorical Transformer-based World Models to continuous action environments for the first time, addressing a significant gap in prior research. To achieve this, we employ a state mixer function that integrates latent state representations with actions, enabling our model to handle continuous control tasks. We validate this approach through qualitative and quantitative analyses on the DeepMind Control Suite, showcasing the effectiveness of Transformer-based World Models in this new domain. Our results highlight the versatility and efficacy of the MaskGIT dynamics prior, paving the way for more accurate world models and effective RL policies. Here you can find our GitHub repository: https://github.com/Cmeo97/GIT-STORM/.
@@ 2.2 Masked Modelling for Visual Representations and Generation
[…]
@@ 3 Method
Following DreamerV3 (Hafner et al., 2023) and STORM (Zhang et al., 2023), we define our framework as a partially observable Markov decision process (POMDP) with discrete timesteps, $t\in\mathbb{N}$, scalar rewards, $r_{t}\in\mathbb{R}$, image observations, $o_{t}\in\mathbb{R}^{h\times w\times c}$, and discrete actions. $a_{t}\in\{1,\ldots,m_{a}\}$. These actions are governed by a policy, $a_{t}\sim\pi(a_{t}\mid o_{1:t},a_{1:t-1})$, where $o_{1:t}$ and $a_{1:t-1}$ represent the previous observations and actions up to timesteps $t$ and t1t-1, respectively. The termination of each episode is represented by a Boolean variable, $c_{t}\in\{0,1\}$. The goal is to learn an optimal policy, $\pi$, that maximizes the expected total discounted rewards, $\mathbb{E}_{\pi}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r_{t}\right]$, where $\gamma\in[0,1]$ serves as the discount factor. The learning process involves two parallel iterative phases: learning the observation and dynamics modules (World Model) and optimizing the policy (Agent).
@@ 3.2 Dynamics Prior Head: MaskGIT Prior
$\mathcal{L_{\mathrm{dyn}}}(\phi)$ $\doteq\max\bigl(1,\operatorname{KL}\!\big[\operatorname{sg}(q_{\phi}(z_{t}\;|\;x_{t}))\circ m_{t}\;\big\|\;\hskip 13.77771ptp_{\phi}(z_{t}\;|\;h_{t-1})\hphantom{)}\big]\bigr)$ (3)
$\mathcal{L_{\mathrm{rep}}}(\phi)$ $\doteq\max\bigl(1,\operatorname{KL}\!\big[\hskip 13.77771ptq_{\phi}(z_{t}\;|\;x_{t})\hphantom{)}\circ m_{t}\;\big\|\;\operatorname{sg}(p_{\phi}(z_{t}\;|\;h_{t-1}))\big]\bigr)$ (4)
@@
[…]
@@ 4.2 Results on Discrete Action Environments: Atari 100k
Figure 3 (Left) compares PI against the baselines. Noticeably, GIT-STORM presents $PI>0.5$ for all baselines, which indicates that, from a probabilistic perspective GIT-STORM would outperform each baseline on a random task Y from Atari 100k with a probability greater than $0.5$. Figure 6 illustrates the Optimality Gap, while Figure 7 presents the fraction of runs with score $\tau$ for different human normalized scores; both confirm the trends observed so far. Moreover, a closer look to Table 5 reveals that GIT-STORM presents an optimality gap of $0.500$, marginally beating DreamerV3, which reports $0.503$ and significantly outperforming all other baselines.
@@ A.1 Observation Module
$\text{Observation encoder: }z_{t}\sim q_{\phi}(z_{t}\mid o_{t})$ (5)
$\text{Observation decoder: }\hat{o}_{t}=p_{\phi}(z_{t})$ (6)
@@ A.2 Policy Learning
$\text{Actor: }a_{t}\sim\pi_{\theta}(a_{t}|s_{t})\ ,\ \text{Critic: }V_{\psi}(s_{t})\approx\mathbb{E}_{\pi_{\theta},p_{\phi}}\left[\sum_{\tau=0}^{\infty}\gamma^{\tau}r_{t+\tau}\right],$ (8)
@@
$\mathcal{L}_{\theta}=\frac{1}{BL}\sum_{n=1}^{B}\sum_{l=1}^{L}\left[-\text{sg}\left(\frac{G_{l}^{\lambda}-V_{\psi}(s_{l})}{\max(1,S)}\right)\ln\pi_{\theta}(a_{l}|s_{l})-\eta H(\pi_{\theta}(a_{l}|s_{l}))\right],\\$ (12)
@@ A.3 Draft and Revise decoding scheme
[…]
@@
4:  $\text{MaskGIT\_codes}\leftarrow\operatorname{MaskGIT\_Codebook}(\bm{z}\circ\mathbf{m}^{i})$
@@
6:  $\text{logits}\leftarrow\xi\odot\text{MaskGIT\_embeddings}$
@@ Appendix C Full Results on RL Task
Table 5: Evaluation on the $26$ games in the Atari 100100k benchmark. We report mean scores as well as aggregated human normalized mean and median, Interquantile Mean (IQM), and Optimality Gap. Following the conventions of Hafner et al. (2021), scores that are the highest or within $5\%$ of the highest score are highlighted in bold.
@@ Appendix D Training Curves
Figure 8: Training profiles across all the checkpoints for the Atari 100k benchmark. The solid line represents the average over 55 seeds while the fill area is defined in terms of maximum and minimum values corresponding to each checkpoint.
Figure 9: Performance evaluation across all the checkpoints for the DMC benchmark for GIT-STORM and STORM. The solid line represents the average over 55 seeds while the fill area is defined in terms of standard deviation values corresponding to each checkpoint.
@@ Appendix H Hyperparameters
Table 11: Action mixer $\zeta_{t}=g_{\theta}(z_{t},a_{t})$. Concatenate denotes combining the last dimension of two tensors and merging them into one new tensor. The variable AA represents the action dimension, which ranges from $3$ to $18$ across different games. $D$ denotes the feature dimension of the Transformer. LN is an abbreviation for layer normalization (Ba et al., 2016).
@@
Random sample ($z_{t}$), Action ($a_{t}$) $32\times 32$, AA